40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9872     Entry time: Mon Apr 28 23:05:03 2014
Author: Jenne 
Type: Update 
Category: LSC 
Subject: ALS CARM and DARM settings 

[Jenne, Koji]

The IFO is being uncooperative tonight, and I have an early morning meeting, so I'm calling it a night. 

Koji's filter module changes have been propagated from the Xarm to the Yarm, to CARM and to DARM.  (Actually, Q overwrote the changes to Xarm on Sunday accidentally, so first he reverted those for us, and then we propagated the changes). 

Today, with careful measuring, we find that for X and Y arms individually locked with the ALS, we want the gains to be +17 for the Yarm, and -17 for the Xarm (with the beatnote up-is-up convention).  This puts the UGFs at 150 Hz. 

We then switched over to CARM and DARM locking.  We guessed that the gains should be a factor of 2 lower since we're pushing on both ETMs for DARM, and the MC2 actuator is roughly the same strength as the sum of the ETMs.  In the end, after measuring the CARM and DARM loops, we find that the gains should be +7.5 for CARM, and +8.0 for DARM to set the UGFs at 150 Hz.  The servo is a little bit delicate, so having too low of gain is not okay. 

For some reason, we seem to be utilizing more actuator range with the new setup, so the limiters in the filter banks have been set to 11,000 (previously were 8,000), and the ALS watch script (ALSdown.py) threshold has been increased to 10,000 (previously 7,000). 

When finding the IR resonances with the new scheme, we are having trouble holding lock throughout the scan.  I have set the tramp for the coarse part of the scan to be 0.05 seconds (previously 0.01 seconds), which is an increase of a factor of 5 in the ramp time.  This helps, but may still not be enough, since we don't always hold lock until both IR resonances are found.

Probably the most annoying thing from tonight is the fact that ETMY keeps drifting off, particularly in yaw, when locked.  I don't have an explanation of why this is happening, but you can watch it happen sometimes, and the lock will be lost shortly thereafter.  Definitely when we lose lock and the ETM gets kicked, it is far enough away in yaw alignment that I have to completely redo the Yarm alignment.  This happens whether or not the ETMY oplevs are on.

To summarize, 3 scripts have been modified:

(1) ALSdown - threshold increased  (Modification from last week - turns off the slow temp servos for the end lasers, clears histories)

(2) ALSfindIRresonance - increase ramp time

(3) Lock_ALS_CARMandDARM - final gain values set to 7.5 for CARM and 8 for DARM, no filters come on until gains all the way up, turns on new set of Koji filters. (Modification from last week - turns on the slow temperature servos for the end lasers)

ELOG V3.1.3-